help
Find the value of y in the equation 3/y-2=8

Answer by stanbon(75887)

Find the value of y in the equation 3/(y-2)=8
Multiply both sides by (y-2) and divide by 8
to get:
y-2=3/8
y=2 3/8 = 19/8
Cheers,
Stan H.